# DualSpeechLM: Towards Unified Speech Understanding and Generation via Dual Speech Token Modeling with Large Language Models

### v1

@@ Appendix A Appendix
Appendix B A Convergence Speed Comparison
@@
Appendix C B Model Details
@@
The AcousticGPT module is implemented as a GPT-style structure, which autoregressively predicts next acoustic tokens based on previous acoustic tokens and conditioned on semantic representations and speaker embeddings. As shown in Figure 5, the module consists of six stacked causal blocks, each comprising a causal multi-head self-attention layer followed by a multi-layer perceptron (MLP). Within each block, layer normalization is applied before both the self-attention and MLP modules, and residual connections are added after each to facilitate gradient flow and improve training stability. The MLP component includes a linear layer, GELU activation, another linear projection, and dropout for regularization. A final layer normalization layer is applied after the last causal block. During training, the hidden states extracted from text LLMs are first selected using the CoC strategy and then concatenated with speaker embeddings extracted by a pre-trained 3D-Speaker encoder (Chen et al. 2025) to form the input to the AcousticGPT module. Then the model generates acoustic tokens that are coherent in both content and speaker characteristics. This architecture enables AcousticGPT to produce high-fidelity speech representations aligned with semantic and stylistic information.
Appendix D C DataSet Statistics
@@
ASR Automatic Speech Recognition Librispeech (Panayotov et al. 2015) 960 280K
S2TT Speech-to-Text Translation CoVoST2-En2De (Wang, Wu, and Pino 2020) 430 290K
@@
SQA Speech Question Answering SQA (Tang et al. 2023) 960 280K
@@
TTS Text-to-Speech LibriTTS-R (Koizumi et al. 2023) 960 350K
T2ST Text-to-Speech Translation CVSS (Jia et al. 2022)-Es2En 60 70K
CVSS (Jia et al. 2022)-Fr2En 109 130K
VC Voice Conversion Libriheavy (Kang et al. 2023) 1K 290K
SC Speech Conversation Synthetic SQA (Tang et al. 2023) 100 28K
@@
SQA Speech Question Answering SQA (Tang et al. 2023) BLEU4/ChatGPTScore
@@
VC Voice Conversion VCTK (Yamagishi et al. 2019) SIM/WER/DNSMOS
SC Speech Conversation Synthetic SQA (Tang et al. 2023) BLEU4/ChatGPTScore
@@
During the training of USTokenizer, we utilize a dataset comprising nearly 2,000 hours of speech, jointly optimizing for multiple spoken language understanding tasks, including Automatic Speech Recognition (ASR), Speech Emotion Recognition (SER), and Speech Question Answering (SQA). Specifically, the LibriSpeech (Panayotov et al. 2015) train set was used for ASR, while the IEMOCAP (Busso et al. 2008) Sessions 1–4 are employed for SER. For the SQA task, we utilize a dataset from SALMONN (Tang et al. 2023), where questions and answers are automatically generated from the LibriSpeech transcripts using ChatGPT. In this setting, the LLM is required to generate text responses to prompts by jointly considering both the speech input and the corresponding text question. We refer to this task as Speech Question Answering (SQA).
@@
As shown in Table 8, in DualSpeechLM, we jointly train a unified model for both speech understanding and generation across eight distinct tasks, leveraging approximately 4,500 hours of speech data. For speech understanding, we use the LibriSpeech (Panayotov et al. 2015) dataset to train and evaluate the Automatic Speech Recognition (ASR) task; the English-to-German (En2De) split of CoVoST2 (Wang, Wu, and Pino 2020) for training and evaluating the Speech-to-Text Translation (S2TT) task; IEMOCAP (Busso et al. 2008) Sessions 1–4 for training and Session 5 for evaluating the Speech Emotion Recognition (SER) task; Similarly, in the SQA task, our DualSpeechLLM is required to generate text responses to prompts by jointly considering both the speech input and the corresponding text question. We use the SQA dataset (Tang et al. 2023), constructed from LibriSpeech using ChatGPT, for both training and evaluation. Specifically, 400 data pairs are randomly selected from the dataset to form a separate test set, which does not overlap with the training set. For speech generation, we employ the LibriTTS-R (Koizumi et al. 2023) dataset for training and evaluating the Text-to-Speech (TTS) task; the Spanish-to-English (Es–En) and French-to-English (Fr–En) splits of  (Jia et al. 2022) for training and evaluating the Text-to-Speech Translation (T2ST) task. For the Voice Conversion (VC) task, we randomly select 1,000 hours from LibriHeavy (Kang et al. 2023) for training. To evaluate the VC task, we randomly construct 400 data pairs from VCTK (Yamagishi et al. 2019) by pairing utterances of the same sentence spoken by different speakers, as well as different utterances by the same speaker for speaker embeddings. Furthermore, we randomly sample 100 hours of speech from the SQA training set and use the open-source Volcengine TTS API to synthesize both corresponding questions and answers into speech, forming speech-question–speech-answer pairs for the Speech Conversation (SC) task. For evaluation, we construct a separate SC test set by randomly selecting 400 additional data pairs and applying the same synthesis procedure. In the SC task, the DualSpeechLLM is required to generate speech responses to speech prompts that include speech input and the corresponding speech question.
@@
As shown in Table 9, we adopt a set of objective metrics to evaluate both understanding and generation performance in DualSpeechLM. Specifically, Word Error Rate (WER) is used in ASR, TTS, and VC to quantify transcription errors by calculating the proportion of substitutions, insertions, and deletions relative to the total number of words in the reference, where lower WER indicates higher transcription fidelity. For TTS and VC tasks, we first transcribe the generated speech using Whisper large-v3 (Radford et al. 2023)  [^3] https://huggingface.co/openai/whisper-large-v3 and then compute WER based on the transcribed output. BLEU4 (Papineni et al. 2002) is employed in S2TT, T2ST, SQA, and SC to measure the n-gram overlap (up to 4-grams) between generated and reference texts, where higher scores reflect stronger textual alignment, especially useful for translation. For SQA and SC, since the answers are typically generated based on the given speech input, BLEU4 can be used to evaluate the consistency between the generated and reference answers. However, as BLEU4 has limited coverage for diverse but correct expressions, we additionally adopt ChatGPTScore for further evaluation for SQA and SC tasks. Specifically, we use ChatGPT 4.1 (Achiam et al. 2023) to directly rate the relevance, fluency, and correctness of generated and reference answers, with the evaluation prompt detailed as shown in Listing 1. For the AC task, we first use Whisper large-v3 (Radford et al. 2023) to transcribe speech answers into text, and then employ the same evaluation prompt with ChatGPT 4.1 to assess the results. For the SER task, we use classification accuracy (ACC) to assess the proportion of correctly predicted emotion categories. For speaker similarity (SIM), used in TTS and VC, we employ the wavlm-base-plus-sv (Chen et al. 2022)  [^4] https://huggingface.co/microsoft/wavlm-base-plus-sv model to extract speaker embeddings from both the prompt and synthesized speech, and compute the cosine similarity between them to assess speaker similarity. Finally, DNSMOS (Reddy, Gopal, and Cutler 2021) is applied in TTS and VC to assess speech naturalness. It is a no-reference neural metric that predicts mean opinion scores (MOS) on a 5-point Likert scale, approximating human judgments of audio quality.
@@
Firstly, you will be given a list of files, which include:
@@
Appendix E D Model Configuration
@@
To verify the generalization of USToken and the effectiveness of the DualSpeechLM framework on different backbones, we conduct experiments using text LLMs of Vicuna-7B [^6] https://huggingface.co/lmsys/vicuna-7b-v1.1 and Phi-3B [^7] https://huggingface.co/microsoft/Phi-3.5-mini-instruct (Abdin et al. 2024). For AcousticGPT, we adopt 6 multi-head self-attention layers, each featuring an embedding dimension of 1024 and 8 attention heads. Causal masking is applied in each layer to effectively model sequential dependencies. During training, we set the weighting hyperparameters both $\lambda$ and $\xi$ in $\mathcal{L}_{\text{generation}}$ to 1. All DualSpeechLM models are trained on 4 NVIDIA A100-40GB GPUs for 60,000 steps. The detailed model and optimization configurations are summarized in Table 11. Furthermore, it is important to note that in SpeechGPT (Zhang et al. 2023a) (Table 1 and Table 2), certain tasks, such as S2TT and T2ST, are not inherently supported. During inference, the S2TT task is accomplished by first applying ASR to transcribe speech into text, followed by translating the resulting English text into German. For the T2ST task, we first translate the Spanish and French text inputs into English, and subsequently perform TTS to synthesize speech.
Appendix F E Discussion

### v3

@@ Appendix A Appendix
Appendix B A. Convergence Speed Comparison
@@
Appendix C B. Model Details
@@
The AcousticGPT module is implemented as a GPT-style structure, which autoregressively predicts next acoustic tokens based on previous acoustic tokens and conditioned on semantic representations and speaker embeddings. As shown in Figure 5, the module consists of six stacked causal blocks, each comprising a causal multi-head self-attention layer followed by a multi-layer perceptron (MLP). Within each block, layer normalization is applied before both the self-attention and MLP modules, and residual connections are added after each to facilitate gradient flow and improve training stability. The MLP component includes a linear layer, GELU activation, another linear projection, and dropout for regularization. A final layer normalization layer is applied after the last causal block. During training, the hidden states extracted from text LLMs are first selected using the CoC strategy and then concatenated with speaker embeddings extracted by a pre-trained 3D-Speaker encoder (chen20243d) to form the input to the AcousticGPT module. Then the model generates acoustic tokens that are coherent in both content and speaker characteristics. This architecture enables AcousticGPT to produce high-fidelity speech representations aligned with semantic and stylistic information.
Appendix D C. DataSet Statistics
@@
ASR Automatic Speech Recognition Librispeech (7178964) 960 280K
S2TT Speech-to-Text Translation CoVoST2-En2De (wang2020covost) 430 290K
@@
SQA Speech Question Answering SQA (tang2023salmonn) 960 280K
@@
TTS Text-to-Speech LibriTTS-R (koizumi2023librittsr) 960 350K
T2ST Text-to-Speech Translation CVSS (jia2022cvss)-Es2En 60 70K
CVSS (jia2022cvss)-Fr2En 109 130K
VC Voice Conversion Libriheavy (kang2023libriheavy) 1K 290K
SC Speech Conversation Synthetic SQA (tang2023salmonn) 100 28K
@@
SQA Speech Question Answering SQA (tang2023salmonn) BLEU4/ChatGPTScore
@@
VC Voice Conversion VCTK (Yamagishi2019CSTRVC) SIM/WER/DNSMOS
SC Speech Conversation Synthetic SQA (tang2023salmonn) BLEU4/ChatGPTScore
@@
During the training of USTokenizer, we utilize a dataset comprising nearly 2,000 hours of speech, jointly optimizing for multiple spoken language understanding tasks, including Automatic Speech Recognition (ASR), Speech Emotion Recognition (SER), and Speech Question Answering (SQA). Specifically, the LibriSpeech (7178964) train set was used for ASR, while the IEMOCAP (busso2008iemocap) Sessions 1–4 are employed for SER. For the SQA task, we utilize a dataset from SALMONN (tang2023salmonn), where questions and answers are automatically generated from the LibriSpeech transcripts using ChatGPT. In this setting, the LLM is required to generate text responses to prompts by jointly considering both the speech input and the corresponding text question. We refer to this task as Speech Question Answering (SQA).
@@
As shown in Table 8, in DualSpeechLM, we jointly train a unified model for both speech understanding and generation across eight distinct tasks, leveraging approximately 4,500 hours of speech data. For speech understanding, we use the LibriSpeech (7178964) dataset to train and evaluate the Automatic Speech Recognition (ASR) task; the English-to-German (En2De) split of CoVoST2 (wang2020covost) for training and evaluating the Speech-to-Text Translation (S2TT) task; IEMOCAP (busso2008iemocap) Sessions 1–4 for training and Session 5 for evaluating the Speech Emotion Recognition (SER) task; Similarly, in the SQA task, our DualSpeechLLM is required to generate text responses to prompts by jointly considering both the speech input and the corresponding text question. We use the SQA dataset (tang2023salmonn), constructed from LibriSpeech using ChatGPT, for both training and evaluation. Specifically, 400 data pairs are randomly selected from the dataset to form a separate test set, which does not overlap with the training set. For speech generation, we employ the LibriTTS-R (koizumi2023librittsr) dataset for training and evaluating the Text-to-Speech (TTS) task; the Spanish-to-English (Es–En) and French-to-English (Fr–En) splits of  (jia2022cvss) for training and evaluating the Text-to-Speech Translation (T2ST) task. For the Voice Conversion (VC) task, we randomly select 1,000 hours from LibriHeavy (kang2023libriheavy) for training. To evaluate the VC task, we randomly construct 400 data pairs from VCTK (Yamagishi2019CSTRVC) by pairing utterances of the same sentence spoken by different speakers, as well as different utterances by the same speaker for speaker embeddings. Furthermore, we randomly sample 100 hours of speech from the SQA training set and use the open-source Volcengine TTS API to synthesize both corresponding questions and answers into speech, forming speech-question–speech-answer pairs for the Speech Conversation (SC) task. For evaluation, we construct a separate SC test set by randomly selecting 400 additional data pairs and applying the same synthesis procedure. In the SC task, the DualSpeechLLM is required to generate speech responses to speech prompts that include speech input and the corresponding speech question.
@@
As shown in Table 9, we adopt a set of objective metrics to evaluate both understanding and generation performance in DualSpeechLM. Specifically, Word Error Rate (WER) is used in ASR, TTS, and VC to quantify transcription errors by calculating the proportion of substitutions, insertions, and deletions relative to the total number of words in the reference, where lower WER indicates higher transcription fidelity. For TTS and VC tasks, we first transcribe the generated speech using Whisper large-v3 (radford2023robustwhisper)  [^3] https://huggingface.co/openai/whisper-large-v3 and then compute WER based on the transcribed output. BLEU4 (papineni2002bleu) is employed in S2TT, T2ST, SQA, and SC to measure the n-gram overlap (up to 4-grams) between generated and reference texts, where higher scores reflect stronger textual alignment, especially useful for translation. For SQA and SC, since the answers are typically generated based on the given speech input, BLEU4 can be used to evaluate the consistency between the generated and reference answers. However, as BLEU4 has limited coverage for diverse but correct expressions, we additionally adopt ChatGPTScore for further evaluation for SQA and SC tasks. Specifically, we use ChatGPT 4.1 (achiam2023gpt) to directly rate the relevance, fluency, and correctness of generated and reference answers, with the evaluation prompt detailed as shown in Figure 6. For the AC task, we first use Whisper large-v3 (radford2023robustwhisper) to transcribe speech answers into text, and then employ the same evaluation prompt with ChatGPT 4.1 to assess the results. For the SER task, we use classification accuracy (ACC) to assess the proportion of correctly predicted emotion categories. For speaker similarity (SIM), used in TTS and VC, we employ the wavlm-base-plus-sv (chen2022wavlm)  [^4] https://huggingface.co/microsoft/wavlm-base-plus-sv model to extract speaker embeddings from both the prompt and synthesized speech, and compute the cosine similarity between them to assess speaker similarity. Finally, DNSMOS (dnsmos) is applied in TTS and VC to assess speech naturalness. It is a no-reference neural metric that predicts mean opinion scores (MOS) on a 5-point Likert scale, approximating human judgments of audio quality.
@@
Firstly, you will be given the following content:
@@
Figure 6: Prompt for QA Evaluation using ChatGPTScore.
Appendix E D. Model Configuration
@@
To verify the generalization of USToken and the effectiveness of the DualSpeechLM framework on different backbones, we conduct experiments using text LLMs of Vicuna-7B [^6] https://huggingface.co/lmsys/vicuna-7b-v1.1 and Phi-3B [^7] https://huggingface.co/microsoft/Phi-3.5-mini-instruct (abdin2024phi3technicalreporthighly). For AcousticGPT, we adopt 6 multi-head self-attention layers, each featuring an embedding dimension of 1024 and 8 attention heads. Causal masking is applied in each layer to effectively model sequential dependencies. During training, we set the weighting hyperparameters both $\lambda$ and $\xi$ in $\mathcal{L}_{\text{generation}}$ to 1. All DualSpeechLM models are trained on 4 NVIDIA A100-40GB GPUs for 60,000 steps. The detailed model and optimization configurations are summarized in Table 11. Furthermore, it is important to note that in SpeechGPT (zhang2023speechgpt) (Table 1 and Table 2), certain tasks, such as S2TT and T2ST, are not inherently supported. During inference, the S2TT task is accomplished by first applying ASR to transcribe speech into text, followed by translating the resulting English text into German. For the T2ST task, we first translate the Spanish and French text inputs into English, and subsequently perform TTS to synthesize speech.
Appendix F E. Semantic Comparison of USTokenizer
Task Prompt
ASR Recognize the speech and give me the transcription.
TTS Please read this sentence out loud.
VC Without altering the spoken content, transform the speaker’s voice in this speech to match the target voice.
T2ST Please translate the [sourcelanguage][source\;language] text into [targetlanguage][target\;language] speech.
SC Please listen to the speech content and provide a spoken answer to the question.
SQA Based on the content, provide a text-based answer to the question.
S2TT Please translate the [sourcelanguage][source\;language] speech into [targetlanguage][target\;language] text transcription.
SER Please describe the emotion of the speaker.
Table 12: Task prompts used in multi-task evaluation.
Evaluation with Recent Tokenizers
To further validate the semantic capability of our USTokenizer, we retrain automatic speech recognition (ASR) models using several recent tokenizers on Librispeech, including TAAE (parkerscaling), MimiCodec (defossez2024moshi), and CosyVoice-2 (du2024cosyvoice2), under identical experimental setups. For fairness, we only use the first quantization layer of MimiCodec, which is distilled from the first layer of WavLM (chen2022wavlm). As shown in Table 13, USTokenizer achieves the lowest Word Error Rate (WER) on both clean and noisy speech conditions. These results demonstrate that USTokenizer provides a more effective abstraction of high-level information, indicating its stronger ability in semantic preservation.
Model Clean Other
$wer\downarrow$
TAAE 13.24 29.49
MimiCodec 5.96 14.54
CosyVoice2 6.29 15.19
USTokenizer(Ours) 4.22 9.71
Table 13: The Comparison with Recent Tokenizers.
Comparison with Baichuan-Audio Tokenizer
We also compare USTokenizer with Baichuan-Audio (li2025baichuanaudio), a recent unified framework for end-to-end speech interaction. Baichuan-Audio adopts an 8-layer Residual Vector Quantization (RVQ) scheme optimized for Mel-spectrogram reconstruction. In contrast, USTokenizer employs a single-layer VQ optimized for self-supervised representation reconstruction, which is simpler, more LLM-friendly, and emphasizes semantic preservation. As summarized in Table 14, Baichuan-Audio exhibits a higher WER, possibly because Mel reconstruction prioritizes acoustic fidelity, which in turn diminishes semantic retention to some extent. By contrast, USTokenizer reconstructs self-supervised representations, achieving a lower bitrate and richer semantic encoding.
Model Frame Rate Token Bitrate Clean/Other
HzHz $num/s$ kbpskbps $wer\downarrow$
Baichuan-Audio 12.5 100 1.08 4.53/10.65
USTokenizer(Ours) 25 25 0.25 4.22/9.71
Table 14: The Comparison with Baichuan-Audio on Librispeech.
Appendix G F. Computational Overhead Analysis
To quantify the additional computational cost introduced by USTokenizer and DualSpeechLM, we report the relative changes in memory and training time under different configurations. As shown in Table 15, the text-LLM module in USTokenizer introduces moderate overhead but remains manageable, as it is unused during inference and converges quickly when applied to DualSpeechLM. Furthermore, DualSpeechLM demonstrates higher training efficiency owing to the low-bitrate token design of USTokenizer, which reduces both memory and time consumption compared with baselines.
Model Setting Memory(%) Time(%)
USTokenizer vs w/o Text LLM $\uparrow$ 288 $\uparrow$ 76
DualSpeechLM vs Baseline-Semantic $\downarrow$ 26 $\downarrow$ 32
DualSpeechLM vs Baseline-Acoustic $\downarrow$ 17 $\downarrow$ 31
Table 15: Training overhead of USTokenizer and DualSpeechLM.
Overall, these results indicate that the additional text-LLM module in USTokenizer slightly increases training cost but does not affect inference efficiency, while DualSpeechLM achieves better computational efficiency.
Appendix H G. Prompts for different tasks
Table 12 summarizes the prompts used for different tasks in training and evaluation of our DualSpeechLM.
Appendix I H. Discussion